# thanks: ZY and HC contributed equally to this workthanks: ZY and HC contributed equally to this work

### v1

@@ I Introduction
The idea that a fundamental pseudoscalar could constitute dark matter (DM) has its origins in the Peccei–Quinn (PQ) solution to the strong-CP problem in QCD [1, 2]. The spontaneous breaking of this symmetry gives rise to a pseudo-Nambu–Goldstone boson known as the QCD axion [3, 4]. Notably, the QCD axion not only solves the strong-CP problem but also emerges as a compelling DM candidate, provided it is produced through non-thermal mechanisms, such as vacuum misalignment, that allow it to remain cold and non-relativistic well before the epoch of recombination [5, 6, 7] (for reviews, see Refs. [8, 9, 10, 11]). While thermal production via interactions with the Standard Model plasma is also possible, it typically leads to hot or warm relics and is therefore only viable for sufficiently heavy and weakly interacting axion-like particles or axions [12, 13, 14]. Such heavy axions that are abundantly produced by thermal processes are, however, unstable to decay into photons. Thus, thermally produced axions are subject to strong constraints on their parameter space [13, 15, 16, 17, 18, 19].
@@
Using approximate formulae and rough comparisons to astrophysical limits, it was observed in Ref. [31] that freeze-in production followed by decay of axions, constrained by the CMB optical depth and other observables [17, 35], seems to place string theory in tension with models of high-temperature reheating. It is the purpose of the present work to place this observation on firm footing with rigorous calculation. Our methodology is the following:
@@ II.1 Axions in string theory
Under these model-building assumptions, the index $i$ in Eq. (2) runs from 11 11 to $h^{1,1}$, and the fields $a_{i}(x)$ give rise to all $h^{1,1}$ axions in the closed-string sector. Dimensionally reducing the 10D action leads to a 4D action for the axions:
@@
To assess the effect of CY topology, we sample 5050 5050 polytopes and 1010 1010 triangulations of each for $h^{1,1}=20,50,100$. For every CY thus constructed, we select every suitable prime toric divisor as a candidate to host QCD, as well as a random intersecting divisor for the electroweak sector. The point in moduli space is specified by Kähler parameters $\vec{t}$, which are given by $\lambda\vec{t}_{0}$, where $\vec{t}_{0}$ is the tip of the stretched Kähler cone and $\lambda$ is chosen such that the QCD divisor takes volume 40 in string units. For the Kähler cone, we use $\mathcal{K}_{\cup}$, which approximates the true CY Kähler cone as a union of toric Kähler cones (see e.g. Ref. [29]). This improves on the calculation of Ref. [31], which used $\mathcal{K}_{V}$, the Kähler cone of the ambient variety. Since $\mathcal{K}_{V}\subset\mathcal{K}_{\cup}$, using the larger $\mathcal{K}_{\cup}$ allows one to include more small divisors. $\mathcal{K}_{\cup}$ is a better approximation to the true Kähler cone of the CY, and the additional small divisors available compared to $\mathcal{K}_{V}$ allow a more representative sampling of the moduli space. Computational constraints limit using $\mathcal{K}_{\cup}$ to smaller values of $h^{1,1}\lesssim 100$, and so we could not explore the largest value $h^{1,1}=491$ as in Ref. [31].
@@ II.2 Axion production and decay
×$\displaystyle\times\left(\frac{a(T_{\mathrm{reh}})}{a(T)}\right)^{3}\left(\frac{m_{a}}{1\,\mathrm{MeV}}\right)\left(\frac{n_{\mathrm{eq}}(T_{\mathrm{reh}})}{1\,\mathrm{MeV}^{3}}\right)\,\mathrm{MeV}^{4}\,,$
@@
×$\displaystyle\times\left(\frac{a(T_{\mathrm{reh}})}{a(T)}\right)^{3}\left(\frac{m_{a}}{1\,\mathrm{MeV}}\right)\,\mathrm{MeV}^{4}\,,$
@@
×$\displaystyle\times\left(\frac{m_{a}}{1\,\mathrm{MeV}}\right)\,.$
@@
×$\displaystyle\times\!\left(\!\frac{T^{\prime}}{T_{\mathrm{Id}}}\frac{\left(4\left(\frac{T_{\mathrm{Id}}}{T^{\prime}}\right)^{2}\!-\!1\right)^{3/2}}{3^{3/2}}\frac{\coth\!{\left(0.1\frac{T_{\mathrm{Id}}}{T^{\prime}}\right)}}{\coth{(0.1)}}\!\right)_{T^{\prime}=\min[T_{\mathrm{Id}},T_{\mathrm{reh}}]}\!,$ (11)
@@ III Ionization of the intergalactic medium by axion decay
$\displaystyle$== $\displaystyle\frac{C_{H}}{(1+z)H(z)}\!\left(\!-\beta_{H}(T_{\gamma})(1\!-\!X_{e})e^{\frac{-E_{H},2s1s}{T_{\gamma}}}\!\!+\!X_{e}^{2}n_{H}\alpha_{H}(T_{b})\!\right)$
@@
For each energy deposition channel, labeled by “$\mathrm{c}$”, with “ion” for ionization, “exc” for excitation, and “heat” for IGM heating, the energy deposition rate takes the form
$\frac{{\rm d}E}{{\rm d}V\,{\rm d}z}\bigg{|}_{\mathrm{dep,c}}=f_{\mathrm{c}}(z)\,\frac{{\rm d}E_{\mathrm{inj}}}{{\rm d}V\,{\rm d}z}\,,$ (16)
where $f_{\mathrm{c}}(z)$ is the redshift-dependent efficiency for the channel “$\mathrm{c}$”. The efficiency function $f_{\mathrm{c}}(z)$ is obtained using the DarkHistory code [37], and we precompute it for our analysis. To simplify the computation, we adopt the beyond on-the-spot assumption [35], in which energy injected at redshift $z$ is assumed to be deposited instantaneously. However, rather than assuming a constant efficiency, we retain the full redshift-dependent form of $f_{\mathrm{c}}(z)$ as computed numerically.
@@
Figure 2: Single axion case. Ionization fraction in the $\Lambda$CDM scenario (black line), compared to a model with one additional axion component, assuming instantaneous energy injection with the “beyond on-the-spot” assumption and using the full energy deposition efficiency function provided by DarkHistory. The axion, constituting a fraction $\approx 10^{-7}$ of the DM, decays at redshift $z_{\rm decay}<1100$, leading to efficient ionization of the intergalactic medium and a CMB optical depth $\tau_{\rm reio}\approx 1.92$.
@@
Figure 4 illustrates the ionization history $X_{e}(z)$ resulting from a representative string axion model within the $h^{1,1}=50$ ensemble. Details for the computations are similar to those outlined for Fig. 2. The selected string axion model includes two axions within the mass range relevant for contributions to the CMB optical depth: one with mass $m_{a}^{(1)}=8.27\,\mathrm{GeV}$ and photon coupling $g_{a\gamma\gamma}^{(1)}=1.87\times 10^{-20}\,\mathrm{GeV}^{-1}$, and a second with $m_{a}^{(2)}=0.72\,\mathrm{GeV}$ and $g_{a\gamma\gamma}^{(2)}=2.1\times 10^{-20}\,\mathrm{GeV}^{-1}$. These axions decay during the early Universe, injecting energy that modifies the evolution of the free electron fraction before $z\approx 10$. The resulting ionization history departs from the standard $\Lambda$CDM scenario, potentially imprinting on the CMB polarization through the enhanced $\tau_{\mathrm{highz}}$. Results are shown for different values of the reheating temperature, up to the maximally allowed one, $T_{\rm reh}^{\rm MAX}$, defined as the value at which the computed optical depth exceeds the $2\sigma$ exclusion threshold from our $\tau_{\mathrm{highz}}$ analysis in Paper I [34]; see Fig. 5 below.
@@ IV The high-$z$ optical depth
$\begin{split}\tau_{\mathrm{lowz}}&=\int^{z_{c}}_{0}\,\sigma_{T}\,n_{e}(z)\,\frac{\mathrm{d}z}{(1+z)H(z)},\\ \tau_{\mathrm{highz}}&=\int^{z_{\mathrm{max}}}_{z_{c}}\,\sigma_{T}\,n_{e}(z)\frac{\mathrm{d}z}{(1+z)H(z)}\,,\end{split}$ (24)
@@
Figure 5: Normalized posterior distribution, $P/P_{\mathrm{max}}$, for the high–redshift contribution to the optical depth, for the critical redshift $z_{c}=30$ defining the separation in Eq. (24). Results are based on the Planck $-\displaystyle\ell$ EE dataset alone and the analysis in Paper I. This posterior distribution gives the upper limit $\tau_{\mathrm{highz}}<0.108$ at 95% CL, which we use to set limits on axion models.
The novelty presented in Paper I [34] consists in employing a Gaussian process (GP) method to sample random reionization histories $X_{e}(z)$ [34]. Using a Markov Chain Monte Carlo (MCMC) analysis with $\mathcal{O}(20)$ parameters describing the GP, and the Planck $-\displaystyle\ell$ E-mode EE-only SimAll likelihood, [^3] The most direct and robust probe of the reionization epoch is provided by the $-\displaystyle\ell$ E-mode polarization data. we derive limits on $X_{e}(z)$ that extend to high redshifts. Our results yield a constraint $\tau_{\mathrm{lowz}}=0.0483^{+0.0083}_{-0.0080}$ at 68% confidence level (CL), and an upper limit of $\tau_{\mathrm{highz}}<0.108$ at 95% CL, using $z_{c}=30$. Figure 5 shows the normalized posterior distribution for the high-redshift contribution to the optical depth, $\tau_{\mathrm{highz}}$, as derived from the Planck $-\displaystyle\ell$ EE polarization data, which serves as our constraint on axion models. The distributions are largely insensitive to the choice of the critical redshift $z_{c}$, as we have demonstrated in Paper I [34]. The code used to cast the reionization results can be found at: github.com/Cheng-Hanyu/CLASS_reio_gpr.
Our GP method extends previous analyses by the Planck collaboration, which used similar techniques to reconstruct the standard low-$z$ reionization history [59]. Our result for $\tau_{\mathrm{lowz}}$ is consistent with their findings. In addition, the extracted posterior for $\tau_{\mathrm{highz}}$ yields robust, model-independent constraints on exotic energy injection. By comparing theoretical predictions for $X_{e}(z)$ that reproduce the empirical posterior of $\tau_{\mathrm{highz}}$, a wide range of scenarios can be directly tested, as we demonstrate below. [^4] In Paper I [34], we validate this approach with a direct $\chi^{2}$ comparison using the Planck $-\displaystyle\ell$ EE-only SimAll likelihood.
@@ V Single axion results
We now apply our method to compute a new result, so far as we are aware, for a single axion: the maximum reheating temperature $T_{\rm reh}^{\rm MAX}$ allowed across the $(m_{a},g_{a\gamma\gamma})$ parameter space. For a given axion mass and photon coupling, we calculate the high-redshift optical depth $\tau_{\mathrm{highz}}$ as a function of reheating temperature, using the freeze-in abundance as input. Figure 7 shows this dependence for a representative selection of single-axion models, as indicated in the legend. The gray dashed line indicates the $2\sigma$ upper bound on $\tau_{\mathrm{highz}}$ derived in Paper I (and shown in Fig. 5), which sets the exclusion threshold. As $T_{\rm reh}$ increases, the freeze-in abundance grows, resulting in an increase in $\tau_{\mathrm{highz}}$. For each model, we identify the point at which $\tau_{\mathrm{highz}}$ crosses the exclusion threshold: this defines the corresponding maximum reheating temperature $T_{\rm reh}^{\rm MAX}$ allowed by current data. For comparison, the black dashed line shows the value $\tau_{\mathrm{highz}}=0.0407$, obtained from the baseline $\Lambda$CDM scenario with $z_{c}=30$ in Eq. (24).
Figure 7: The optical depth $\tau_{\mathrm{highz}}$ as a function of reheating temperature, for a selection of single-axion models with varying mass and photon coupling. The gray dashed line marks the $2\sigma$ upper limit on $\tau_{\mathrm{highz}}$ derived in Paper I (see also Fig. 5), while the black dashed line shows the $\Lambda$CDM contribution (see main text for details). The intersection of each curve with the gray dashed line defines the maximum reheating temperature $T_{\rm reh}^{\rm MAX}$ allowed for that model.

### v2

@@ I Introduction
The idea that a fundamental pseudoscalar could constitute dark matter (DM) has its origins in the Peccei–Quinn (PQ) solution to the strong-CP problem in QCD [1, 2]. The spontaneous breaking of this symmetry gives rise to a pseudo-Nambu–Goldstone boson known as the QCD axion [3, 4]. Notably, the QCD axion not only solves the strong-CP problem but also emerges as a compelling DM candidate, provided it is produced through non-thermal mechanisms, such as vacuum misalignment, that allow it to remain cold and non-relativistic well before the epoch of recombination [5, 6, 7] (for reviews, see Refs. [8, 9, 10, 11]). While thermal production via interactions with the Standard Model plasma is also possible, it typically leads to hot or warm relics and is therefore only viable for sufficiently heavy and weakly interacting axion-like particles or axions [12, 13, 14]. Such heavy axions that are abundantly produced by thermal processes are, however, unstable to decay into photons. Thus, thermally produced axions are subject to strong constraints on their parameter space [13, 15, 16, 17, 18, 19, 20, 21, 22].
@@
Using approximate formulae and comparisons to astrophysical limits, it was observed in Ref. [34] that freeze-in production followed by decay of axions, constrained by the CMB optical depth and other observables [17, 38], seems to place string theory in tension with models of high-temperature reheating. It is the purpose of the present work to place this observation on firm footing with rigorous calculation. Our methodology is the following:
@@ II.1 Axions in string theory
Under these model-building assumptions, the index $i$ in Eq. (2) runs from $1$ to $h^{1,1}$, and the fields $a_{i}(x)$ give rise to all $h^{1,1}$ axions in the closed-string sector. Dimensionally reducing the 10D action leads to a 4D action for the axions:
@@
To assess the effect of CY topology, we sample $50$ polytopes and $10$ triangulations of each for $h^{1,1}=20,50,100$. For every CY thus constructed, we select every suitable prime toric divisor as a candidate to host QCD, as well as a random intersecting divisor for the electroweak sector. The point in moduli space is specified by Kähler parameters $\vec{t}$, which are given by $\lambda\vec{t}_{0}$, where $\vec{t}_{0}$ is the tip of the stretched Kähler cone and $\lambda$ is chosen such that the QCD divisor takes volume 40 in string units. For the Kähler cone, we use $\mathcal{K}_{\cup}$, which approximates the true CY Kähler cone as a union of toric Kähler cones (see e.g. Ref. [32]). This improves on the calculation of Ref. [34], which used $\mathcal{K}_{V}$, the Kähler cone of the ambient variety. Since $\mathcal{K}_{V}\subset\mathcal{K}_{\cup}$, using the larger $\mathcal{K}_{\cup}$ allows one to include more small divisors. $\mathcal{K}_{\cup}$ is a better approximation to the true Kähler cone of the CY, and the additional small divisors available compared to $\mathcal{K}_{V}$ allow a more representative sampling of the moduli space. Computational constraints limit using $\mathcal{K}_{\cup}$ to smaller values of $h^{1,1}\lesssim 100$, and so we could not explore the largest value $h^{1,1}=491$ as in Ref. [34].
@@ II.2 Axion production and decay
$\displaystyle\times\left(\frac{a(T_{\mathrm{reh}})}{a(T)}\right)^{3}\left(\frac{m_{a}}{1\,\mathrm{MeV}}\right)\left(\frac{n_{\mathrm{eq}}(T_{\mathrm{reh}})}{1\,\mathrm{MeV}^{3}}\right)\,\mathrm{MeV}^{4}\,$,
@@
$\displaystyle\times\left(\frac{a(T_{\mathrm{reh}})}{a(T)}\right)^{3}\left(\frac{m_{a}}{1\,\mathrm{MeV}}\right)\,\mathrm{MeV}^{4}\,$,
@@
$\displaystyle\times\left(\frac{m_{a}}{1\,\mathrm{MeV}}\right)\,.$
@@
$\displaystyle\times\!\left(\!\frac{T^{\prime}}{T_{\mathrm{Id}}}\frac{\left(4\left(\frac{T_{\mathrm{Id}}}{T^{\prime}}\right)^{2}\!-\!1\right)^{3/2}}{3^{3/2}}\frac{\coth\!{\left(0.1\frac{T_{\mathrm{Id}}}{T^{\prime}}\right)}}{\coth{(0.1)}}\!\right)_{T^{\prime}=\min[T_{\mathrm{Id}},T_{\mathrm{reh}}]}\!,$ (11)
@@ III Ionization of the intergalactic medium by axion decay
$\displaystyle=$ $\displaystyle\frac{C_{H}}{(1+z)H(z)}\!\left(\!-\beta_{H}(T_{\gamma})(1\!-\!X_{e})e^{\frac{-E_{H},2s1s}{T_{\gamma}}}\!\!+\!X_{e}^{2}n_{H}\alpha_{H}(T_{b})\!\right)$
@@
For each energy deposition channel, labeled by $\mathrm{c}$”, with “ion” for ionization, “exc” for excitation, and “heat” for IGM heating, the energy deposition rate takes the form
$\frac{{\rm d}E}{{\rm d}V\,{\rm d}z}\bigg|_{\mathrm{dep,c}}\equiv f_{\mathrm{c}}(z)\,\frac{{\rm d}E_{\mathrm{inj}}}{{\rm d}V\,{\rm d}z}\,$, (16)
where $f_{\mathrm{c}}(z)$ is the redshift-dependent efficiency for the channel $\mathrm{c}$”. The efficiency function $f_{\mathrm{c}}(z)$ is obtained using the DarkHistory code [40], and we precompute it for our analysis. To simplify the computation, we adopt the beyond on-the-spot assumption [38], in which energy injected at redshift $z$ is assumed to be deposited instantaneously. However, rather than assuming a constant efficiency, we retain the full redshift-dependent form of $f_{\mathrm{c}}(z)$ as computed numerically.
@@
[…]
@@
[…]
@@ IV The high-$z$ optical depth
[…]
@@
Figure 5: Normalized posterior distribution, $P/P_{\mathrm{max}}$, for the high–redshift contribution to the optical depth, for the critical redshift $z_{c}=30$ defining the separation in Eq. (24). Results are based on the Planck low-$\ell$ $E$ dataset alone and the analysis in Paper I. This posterior distribution gives the upper limit $\tau_{\mathrm{highz}}<0.108$ at 95% CL, which we use to set limits on axion models.
The novelty presented in Paper I [37] consists in employing a Gaussian process (GP) method to sample random reionization histories $X_{e}(z)$ [37]. Using a Markov Chain Monte Carlo (MCMC) analysis with $\mathcal{O}(20)$ parameters describing the GP, and the Planck low-$\ell$ E-mode EE-only SimAll likelihood, [^3] The most direct and robust probe of the reionization epoch is provided by the low-$\ell$ E-mode polarization data. we derive limits on $X_{e}(z)$ that extend to high redshifts. Our results yield a constraint $\tau_{\mathrm{lowz}}=0.0483^{+0.0083}_{-0.0080}$ at 68% confidence level (CL), and an upper limit of $\tau_{\mathrm{highz}}<0.108$ at 95% CL, using $z_{c}=30$. Figure 5 shows the normalized posterior distribution for the high-redshift contribution to the optical depth, $\tau_{\mathrm{highz}}$, as derived from the Planck low-$\ell$ EE polarization data, which serves as our constraint on axion models. The distributions are largely insensitive to the choice of the critical redshift $z_{c}$, as we have demonstrated in Paper I [37]. The code used to cast the reionization results can be found at: github.com/Cheng-Hanyu/CLASS_reio_gpr.
Our GP method extends previous analyses by the Planck collaboration, which used similar techniques to reconstruct the standard low-$z$ reionization history [62]. Our result for $\tau_{\mathrm{lowz}}$ is consistent with their findings. In addition, the extracted posterior for $\tau_{\mathrm{highz}}$ yields robust, model-independent constraints on exotic energy injection. By comparing theoretical predictions for $X_{e}(z)$ that reproduce the empirical posterior of $\tau_{\mathrm{highz}}$, a wide range of scenarios can be directly tested, as we demonstrate below. [^4] In Paper I [37], we validate this approach with a direct $\chi^{2}$ comparison using the Planck low-$\ell$ EE-only SimAll likelihood.
@@ V Single axion results
[…]